The rate at which a super computer body cools is proportional to the difference between the temperature of the body and that of the surrounding air. If a body in air at 25°C will cool from 100°C to 75°C in one minute, find its temperature at the end of three minutes.


1
Expert's answer
2022-04-17T23:19:45-0400

"\int \frac{dU}{U-25}=k\smallint dt+ln"C


"ln(U-25)=kt+lnC"

"U-25=Ce^{kt}"

For t=0, U=100

For t=1, U=75

"100-25=Ce^{0k}"

C=75

"75-25=75e^{1k}"

"e^k=0.67"

k=-0.4

"U(3)-25=75e^{-0.4x3}"

U(3)=47.6 C